buestion 3 A night club owner has both student and adult customers. The demand for drinks by a typical student is Q. 18-3P. The demand for drinks by a typical adult is QA-10-2P. There are an equal number of adults and students. The marginal cost of each drink is $2. 6 18 Q 10 Part 1: a) What price will the club owner set if she cannot discriminate between the two groups? What will total profit be at this price? b) If the club owner could separate the groups and practice 3rd-degree price discrimination what price per drink would be charged to members of each group? What would be the club owners profit? Part 2: a) If the club owner in problem 4 can card patrons and determine who among them is a student and who is not and in turn can serve each group by offering a cover charge and a number of drink tokens to each group ( 2-part pricing what will the cover charge and number of tokens to students be? what will the cover charge and number of tokens to adults be? What is the club-owners profit under this regime b) Is total welfare higher with the 2-part pricing plan than with 3rd degree price discrimination (in part 1b above)? Explain your answer
Answer #1

The demand of a typical student for drinks is, Of = 18-3P The demand of a typical adult for drinks is, Q+ = 10-2P The marginaMR = MC 28-20 =10 Q=9 II i II P=3.8 Hence, the club owner will charge price of $3.5 when she cannot discriminate. The total pMRⓇ =6=.3 At equilibrium MR = MC O Q* = 6 wla A P=$4 Hence, she will charge $4 per drink to students. Adults The inverse demaMR = MC 5-4 = 2 Q4 = 3 P=s- P=$3.5 Hence, she will charge $3.5 per drink to adults. The profit of club owner will be Profit =Cover charge = -(6-2)(12)+2(12) = 24+24 = 48 The cover charge for adults will be, Cover charge* = =(5-2)6)+2(6) =9+12 = 21 He